### v1

Abstract
Model predictive control (MPC) is widely used in process control due to its interpretability and ability to handle constraints. As a parametric policy in reinforcement learning (RL), MPC offers strong initial performance and low data requirements compared to black-box policies like neural networks. However, most RL methods rely on first-order updates, which scale well to large parameter spaces but converge at most linearly, making them inefficient when each policy update requires solving an optimal control problem, as is the case with MPC. While MPC policies are typically sparsely parameterized and thus amenable to second-order approaches, existing second-order methods demand second-order policy derivatives, which can be computationally and memory-wise intractable.
This work introduces a Gauss-Newton approximation of the deterministic policy Hessian that eliminates the need for second-order policy derivatives, enabling superlinear convergence with minimal computational overhead. To further improve robustness, we propose a momentum-based Hessian averaging scheme for stable training under noisy estimates. We demonstrate the effectiveness of the approach on a nonlinear continuously stirred tank reactor (CSTR), showing faster convergence and improved data efficiency over state-of-the-art first-order methods.
keywords:
[…]
1 Introduction
Model predictive control (MPC) is a popular approach for the control of chemical processes because it can deal with highly dynamic, nonlinear multivariable systems subject to safety and quality constraints. In MPC, a system model of the controlled plant is used to predict and optimize the future behavior of the system states so that a objective function is minimized while ensuring process constraints. The first control action of the optimal solution is then deployed to the real plant, the next state is measured or estimated, and the process is repeated. Since MPC delivers a sequence of optimal control inputs and a sequence of upcoming states, the control policy is highly interpretable. As an established method, many theoretical results regarding stability and feasibility exist (Rawlings et al., 2024).
Besides rigorous model-based optimal control approaches like MPC, there are emerging data-driven approaches such as reinforcement learning (RL) that show suitability and success in the process industry (Yoo et al., 2021; Gopaluni et al., 2020; Lee and Lee, 2007). Model-free RL as a subfield RL focuses on learning an optimal policy through trial-and-error interaction between an agent (controller) and its environment (plant), without relying on a model of the environment. In practice, this optimal policy is often approximated directly using a rich parametric function approximator such as neural networks (NN). Contrary to MPC, this has the benefit that no differentiable system model is necessary to perform online optimization, which is especially interesting if models are hard to build or impossible to differentiate. However, in general model-free RL approaches need large amounts of data, which can be prohibitive for application to industrial scale processes (Recht, 2019).
We focus in this work on two main reasons for the large data requirements of RL: the difficulty of initialization and the use of first-order optimization approaches. If expert knowledge is available beforehand, the NN policy of the RL agent can be pretrained to delivers a reasonable initial performance. However, if expert knowledge is not available, the NN policy is often initialized with random values such that many updates are necessary to achieve a similar performance. Most popular model-free RL algorithms only consider update rules that consist of first-order derivative information such as Fujimoto et al. (2018) and Haarnoja et al. (2018). In either case, the maximum convergence rate of first-order approaches is limited to be linear (Bottou et al., 2018). Methods that consider (approximate) second-order derivative information such as Newton, quasi-Newton, or inexact Newton methods can provide up to quadratic convergence. However, all those approaches must store at least a Hessian matrix approximation in memory and a linear system of equations must be solved for each update step at least approximately. This can be prohibitive especially when rich parameterized policy approximators such as NNs are used. Works such as Korbit et al. (2024), Liu et al. (2024) and Yao et al. (2021) address the memory issue by approximating the Hessian via scalars or diagonal representations, whereas works such as Chauhan et al. (2020) and Schraudolph et al. (2007) only solve the linear system of equations inexactly.
Recent research identifies the challenges described above and proposes to use MPC instead of NNs as a policy for the RL agent (Gros and Zanon, 2020; Banker and Mesbah, 2025; Brandner and Lucia, 2024; Mayfrank et al., 2024). Figure 1 illustrates how NN policies can be replaced by MPC. Reiter et al. (2025) provide a comprehensive review for MPC used as policy in RL. The fundamental idea is to introduce tunable parameters to the MPC’s optimal control problem and to adjust them in the same way as the parameters in NN policies. Contrary to NN policies, MPC provides a reasonable initial policy since the proposed control action results from the solution of a model-based optimal control approach. Further, MPC typically has fewer tuning parameters than NN policies. Due to the sparse parameterization of MPC policies, Hessian computation becomes tractable, allowing more sophisticated updates with superlinear convergence rates, which are of significant interest for MPC policies because each RL transition requires the solution of an optimization problem, which is computationally expensive.
Figure 1: Comparison between deep RL (left) and MPC-based RL (right).
Different update rules were proposed in the past to improve the data efficiency. Kakade (2001) introduces the natural policy gradient, which scales the policy gradient by the inverse of the Fisher information matrix. Although being limited to linear convergence, it often shows overall practical improvement. Further improvement is obtained via quasi-Newton approaches with a superlinear convergence rate as proposed in Furmston et al. (2016), Jha et al. (2020), and Kordabad et al. (2022). With a special focus on MPC policies, Brandner and Lucia (2024) propose a tailored algorithm that reduces the number of updates significantly. However, the second-order policy derivative with respect to its parameters needs to be computed. Brandner and Lucia (2024) show that the second-order derivative of MPC policies can be obtained by the solution of a linear system of equations. Unfortunately, the right-hand side of the linear system involves third-order derivatives of the MPC’s Lagrangian with respect to all primal-dual variables, which can be computationally demanding and memory intensive, especially for high dimensional and highly nonlinear system models.
Lastly, it needs to be said that all considered update rules, regardless if first or second-order, require to compute expected values over the distribution of states that appear under closed-loop operation, and further assume perfect knowledge of the action-value function, which evaluates how good a certain action in a given state is, when following the present policy. In practice however, these expected values can typically only be estimated empirically by a sufficiently large set of closed-loop samples, whereas the action-value function can only be approximated. Both factors contribute to the fact that the gradient and the Hessian cannot be computed perfectly, which can lead to unstable training especially if the Hessian eigenvalues are estimated wrongly.
In this work, we address both the high memory requirements and high computational demand when using MPC policies in second-order RL due to the second-order policy derivatives, and the training instabilities due to poor gradient and Hessian estimation. Our main contributions are twofold:
  1. [1)]
  2. 1.
    We propose a Gauss-Newton approximation of the deterministic policy Hessian that does not require second-order policy derivatives, leading to significant computational savings. We rigorously prove that our approach can still lead to superlinear convergence despite the simpler required computations. Our proposed approximation alleviates the high data requirements compared to the established first-order policy updates in RL.
  3. 2.
    We derive a momentum-based update rule for Hessian estimation that returns an unbiased exponentially averaged Hessian estimate. This estimate is robust with respect to noise in local Hessian samples and therefore enables stable and fast training as it reduces the noise in the eigenvalues of the Hessian. A reliable Hessian estimate is very relevant for MPC policies in RL because the influence of the policy parameters on the objective can typically vary by several orders of magnitude.
The work is structured as follows. We introduce the relevant preliminaries on Markov decision processes, iterative policy optimization via RL, and MPC as policy approximator in RL in Section 2. We present the proposed Gauss-Newton Hessian approximation and momentum-based update rule in Section 3. We illustrate the superlinear convergence capability of the Gauss-Newton approximation on an analytical case study in Section 4 and show the practical improvement of the Gauss-Newton approximation and the proposed momentum-based update rule at the example of a nonlinear continuously stirred tank reactor (CSTR) in Section 5. Section 6 concludes the work and provides perspectives for future work.
2 Preliminaries
2.1 Markov decision processes
RL can solve problems that can be formalized as a Markov decision process (MDP). MDPs are a general mathematical modeling framework for sequential decision making (Sutton and Barto, 2018). In an MDP, an agent (controller) interacts with an environment (plant) as depicted in Figure 1. The agent computes a certain action $\boldsymbol{a}\in\mathcal{A}\subseteq\mathbb{R}^{n_{\boldsymbol{a}}}$ according to the agent’s decision making policy $\pi$, given the current environmental state $\boldsymbol{s}\in\mathcal{S}\subseteq\mathbb{R}^{n_{\boldsymbol{s}}}$. In this work, we focus on deterministic polices $\pi:\mathcal{S}\rightarrow\mathcal{A}$ that directly map the current state $\boldsymbol{s}$ to its proposed action $\boldsymbol{a}$. This action $\boldsymbol{a}$ is handed to the environment, which transitions from the current state $\boldsymbol{s}$ to a subsequent state $\boldsymbol{s}^{+}\in\mathcal{S}$ according to the MDP’s conditional transition probability $\rho:\mathcal{S}\times\mathcal{S}\times\mathcal{A}\rightarrow[0,\infty)$. The subsequent state is sampled from $\rho$ according to
$\boldsymbol{s}^{+}\sim\rho(\cdot|\boldsymbol{s},\boldsymbol{a}).$ (1)
Once that the environment transitioned from state $\boldsymbol{s}$ to $\boldsymbol{s}^{+}$, the new environmental state is handed to the agent together with a reward $r\in\mathbb{R}$. This reward $r$ is computed according the environment’s reward function $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, which depends on the current state $\boldsymbol{s}$ and the taken action $\boldsymbol{a}$, and serves as an immediate measure of how good the current state-action pair $\boldsymbol{s},\boldsymbol{a}\right>$ is, while neglecting future consequences due to upcoming subsequent states.
The general goal under which an MDP is solved, is to find a policy that maximizes not only the immediate reward but the expected future cumulative reward. To measure the expected future cumulative reward in a given state, the concept of value functions are used. The state-value function $V^{\pi}:\mathcal{S}\rightarrow\mathbb{R}$ measures how good policy $\pi$ performs when being in state $\boldsymbol{s}$. It is defined as the expected sum of rewards when being in state $\boldsymbol{s}$ and taking the action proposed by the policy $\boldsymbol{a}=\pi(\boldsymbol{s})$
[…]
$=\mathbb{E}\left[\left.\sum_{i=0}^{\infty}\gamma^{i}R(\boldsymbol{s}_{i},\pi(\boldsymbol{s}_{i}))\right|\boldsymbol{s}_{0}=\boldsymbol{s},\boldsymbol{s}_{i+1}\sim\rho(\cdot|\boldsymbol{s}_{i},\pi(\boldsymbol{s}_{i}))\right]$ (2)
with $0<\gamma\leq 1$ being a discount factor that allows to balance immediate and future reward. All expected values are taken over the conditional transition probability $\rho(\cdot|\boldsymbol{s},\pi(\boldsymbol{s}))$ under closed-loop operation. The state-value function can also written recursively as
$V^{\pi}(\boldsymbol{s})=R(\boldsymbol{s},\pi(\boldsymbol{s}))+\gamma\mathbb{E}_{\boldsymbol{s}^{+}\sim\rho(\cdot|\boldsymbol{s},\pi(\boldsymbol{s}))}\left[V^{\pi}(\boldsymbol{s}^{+})\right].$ (3)
The action-value function $Q^{\pi}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ generalizes the state-value function to taken actions that are different from actions proposed by the policy. Analogous to the state-value function, the action-value function is recursively defined as
$Q^{\pi}(\boldsymbol{s},\boldsymbol{a})=R(\boldsymbol{s},\boldsymbol{a})+\gamma\mathbb{E}_{\boldsymbol{s}^{+}\sim\rho(\cdot|\boldsymbol{s},\boldsymbol{a})}\left[V^{\pi}(\boldsymbol{s}^{+})\right].$ (4)
In general, the policy should be optimal not only for one single state $\boldsymbol{s}$ but for all possible states. To generalize the local state-dependent policy performance measured by $V^{\pi}(\boldsymbol{s})$ to all possible states that can occur under the MDP, the expected cumulative reward $J$ is introduced. It is defined as the expected value of the state-values for all initial states $\boldsymbol{s}_{0}\in\mathcal{S}_{0}\subseteq\mathcal{S}$ that can occur when being sampled from the distribution $\rho_{0}:\mathcal{S}_{0}\rightarrow[0,\infty)$
$J(\pi)=\mathbb{E}_{\boldsymbol{s}_{0}\sim\rho_{0}}\left[V^{\pi}(\boldsymbol{s}_{0})\right].$ (5)
A policy $\pi^{\star}$ is considered optimal (denoted with ) when it maximizes the expected cumulative reward
$\pi^{\star}=\arg\max_{\pi}J(\pi).$ (6)
The following inequalities can be derived directly from the optimality of the policy $\pi^{\star}$
$V^{\pi^{\star}}(\boldsymbol{s})$ $\geq V^{\pi}(\boldsymbol{s}),\quad\forall\boldsymbol{s}\in\mathcal{S}.$ (7)
$Q^{\pi^{\star}}(\boldsymbol{s},\pi^{\star}(\boldsymbol{s}))$ $\geq Q^{\pi^{\star}}(\boldsymbol{s},\boldsymbol{a}),\quad\forall\boldsymbol{s},\boldsymbol{a}\in\mathcal{S}\times\mathcal{A}.$ (8)
2.2 Iterative policy optimization in reinforcement learning
To find the policy $\pi^{\star}$ that solves (6), an infinite dimensional optimization problem over the space of all admissible policy functions must be solved. This is intractable in most cases. Therefore, the policy is often parameterized by a function approximator $\pi_{\boldsymbol{\theta}}(\boldsymbol{s})$, where $\boldsymbol{\theta}\in\mathbb{R}^{n_{\boldsymbol{\theta}}}$ are tunable policy parameters (Sutton and Barto, 2018). The expected cumulative reward transforms from a function of the policy $J(\pi)$ into a function of the policy parameters $J(\pi_{\boldsymbol{\theta}})=J(\boldsymbol{\theta})$. The infinite dimensional optimization problem (6) simplifies to the finite dimensional optimization problem
$\boldsymbol{\theta}^{\star}=\arg\max_{\boldsymbol{\theta}}J(\boldsymbol{\theta}).$ (9)
In its general form, optimization problem (9) can be categorized as an unconstrained nonlinear program with stochastic objective since $J$ is defined in terms of expected values as shown in (5). Iterative optimization algorithms are standard tools to solve such optimization problems numerically (Nocedal and Wright, 2006). The basic idea is to start at an initial guess $\boldsymbol{\theta}_{0}$ of the optimal solution and improve the guess iteratively. The general iterative update process at the $k$-th iteration step can be described as
$\boldsymbol{\theta}_{k+1}=\boldsymbol{\theta}_{k}+\boldsymbol{p}_{k}$ (10)
with $\boldsymbol{p}_{k}\in\mathbb{R}^{n_{\boldsymbol{\theta}}}$ denoting the applied parameter update. The way that the parameter update $\boldsymbol{p}_{k}$ is computed depends on the chosen update rule, which severely influences the performance of the iterative optimization. Different update rules will be presented in the following subsections. The iteration process is repeated until the iterates converge to a point that satisfies the first-order optimality condition for unconstrained optimization sufficiently well
$\left\|\nabla_{\boldsymbol{\theta}}J(\boldsymbol{\theta})\right\|\leq\epsilon$ (11)
[…]
2.2.1 First-order policy optimization
An optimization algorithm is considered to be of first-order, if it only uses first-order derivative information of the optimization problem during the optimization process. In case of (9), this refers to the gradient of the expected cumulative reward with respect to the policy parameters $\nabla_{\boldsymbol{\theta}}J(\boldsymbol{\theta})\in\mathbb{R}^{n_{\boldsymbol{\theta}}}$. In order to preserve a concise and clutter-free notation, we introduce the gradient of the objective function, evaluated at the current iterate $\boldsymbol{\theta}_{k}$ as $\boldsymbol{g}_{k}=\nabla_{\boldsymbol{\theta}}J(\boldsymbol{\theta})|_{\boldsymbol{\theta}=\boldsymbol{\theta}_{k}}$. Silver et al. (2014) derive the deterministic policy gradient as
$\nabla_{\boldsymbol{\theta}}J(\boldsymbol{\theta})=\mathbb{E}_{\boldsymbol{s}}\left[\nabla_{\boldsymbol{\theta}}\pi_{\boldsymbol{\theta}}^{\top}(\boldsymbol{s})\nabla_{\boldsymbol{a}}Q^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s},\boldsymbol{a})|_{\boldsymbol{a}=\pi_{\boldsymbol{\theta}}(\boldsymbol{s})}\right]$ (12)
where $\nabla_{\boldsymbol{\theta}}\pi_{\boldsymbol{\theta}}(\boldsymbol{s})\in\mathbb{R}^{n_{\boldsymbol{a}}\times n_{\boldsymbol{\theta}}}$ denotes the Jacobian of the policy with respect to its parameters and $\nabla_{\boldsymbol{a}}Q^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s},\boldsymbol{a})\in\mathbb{R}^{n_{\boldsymbol{a}}}$ denotes the gradient of the policy’s action-value function with respect to the taken actions, which is evaluated at the policy’s proposed action $\boldsymbol{a}=\pi_{\boldsymbol{\theta}}(\boldsymbol{s})$. The expected value is taken over all $\boldsymbol{s}$ that come up during closed-loop operation.
As the simplest form of a first-order method for stochastic objectives, we introduce the stochastic gradient ascent algorithm (Robbins and Monro, 1951). The stochastic gradient ascent algorithm stems from the deterministic gradient ascent algorithm, which updates the current iterate $\boldsymbol{\theta}_{k}$ into the direction of steepest ascent, which is the gradient of the objective function at the current iterate $\boldsymbol{g}_{k}$. Since the deterministic policy gradient (12) is computed using an expected value, it can only be estimated empirically using a finite amount of samples $N\in\mathbb{N}$ such that
$\tilde{\boldsymbol{g}}_{k}=\frac{1}{N}\sum_{i=1}^{N}\nabla_{\boldsymbol{\theta}}\pi_{\boldsymbol{\theta}}^{\top}(\boldsymbol{s}_{i})\nabla_{\boldsymbol{a}}Q^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s}_{i},\boldsymbol{a})|_{\boldsymbol{a}=\pi_{\boldsymbol{\theta}}(\boldsymbol{s}_{i})}$ (13)
with $\tilde{\boldsymbol{g}}_{k}\approx\boldsymbol{g}_{k}$ leading to noisy estimates of the true deterministic policy gradient. The parameter update $\boldsymbol{p}_{k}$ for the stochastic gradient ascent algorithm is computed as
$\boldsymbol{p}_{k}=\alpha\tilde{\boldsymbol{g}}_{k}$ (14)
where $\alpha>0$ is a scalar step size, also often called the learning rate.
Many works aim to improve the stability and the convergence of stochastic gradient ascent by applying an exponential moving average to observed gradient estimates (Kingma and Ba, 2015; Duchi et al., 2011; Rumelhart et al., 1986). These methods are called momentum-based methods. In its core, momentum-based methods update an exponential moving average $\boldsymbol{m}_{k}\in\mathbb{R}^{n_{\boldsymbol{\theta}}}$ of past gradients with the newly obtained gradient estimate $\tilde{\boldsymbol{g}}_{k}$ according to
$\boldsymbol{m}_{k}=\beta\boldsymbol{m}_{k-1}+(1-\beta)\tilde{\boldsymbol{g}}_{k}$ (15)
where $0\leq\beta\leq 1$ is an exponential decay factor. The larger $\beta$ the slower the moving average $\boldsymbol{m}_{k}$ adapts to the current gradient estimate. Typically, the moving average is initialized as a zero vector $\boldsymbol{m}_{0}=0$. This causes initialization bias of the exponential moving average towards zero, which results in slow improvement in early stages of the optimization. Kingma and Ba (2015) correct the initialization bias by
$\hat{\boldsymbol{m}}_{k}=\frac{\boldsymbol{m}_{k}}{(1-\beta^{k})}$ (16)
with $\hat{\boldsymbol{m}}_{k}\in\mathbb{R}^{n_{\boldsymbol{\theta}}}$ being the bias-corrected exponential moving average. The update rule of a the bias-corrected momentum-based stochastic gradient ascent method is similar to (14) but replaces the gradient with the bias-corrected average
$\boldsymbol{p}_{k}=\alpha\hat{\boldsymbol{m}}_{k}.$ (17)
2.2.2 Second-order policy optimization
Second-order methods improve upon first-order methods by also considering second-order derivative information. The commonly used Newton method derives the following update rule
$\boldsymbol{p}_{k}=-\boldsymbol{H}_{k}^{-1}\boldsymbol{g}_{k}$ (18)
where $\boldsymbol{H}_{k}$ denotes the general Hessian matrix of the objective function with respect to its decision variables at the $k$-th iterate. It can be shown that this update rule delivers a quadratic convergence rate (Nocedal and Wright, 2006). In case of (9), $\boldsymbol{H}_{k}$ refers to the Hessian of the expected cumulative reward with respect to the policy parameters $\nabla_{\boldsymbol{\theta}}^{2}J(\boldsymbol{\theta})\in\mathbb{R}^{n_{\boldsymbol{\theta}}\times n_{\boldsymbol{\theta}}}$ evaluated at the $k$-th iterate $\boldsymbol{H}_{k}=\nabla_{\boldsymbol{\theta}}^{2}J(\boldsymbol{\theta})|_{\boldsymbol{\theta}=\boldsymbol{\theta}_{k}}$. However, the computation of the exact deterministic policy Hessian is computationally expensive and therefore prohibitive for most applications.
Quasi-Newton methods are a computationally tractable alternative to exact Newton methods not only for RL but for optimization in general. The main idea of quasi-Newton methods is to approximate the exact Hessian with a computationally tractable alternative $\boldsymbol{B}_{k}\approx\boldsymbol{H}_{k}$ at the expense of a reduced superlinear convergence rate (Nocedal and Wright, 2006). In alignment to Newton methods, the exact Hessian is interchanged with the tractable approximation resulting in the following update rule for quasi-Newton methods
$\boldsymbol{p}_{k}=-\boldsymbol{B}_{k}^{-1}\boldsymbol{g}_{k}.$ (19)
With a focus on policy optimization via RL, Kordabad et al. (2022) propose a computationally cheaper approximation of the deterministic policy Hessian. The authors show that the following expression approximates the deterministic policy Hessian $\boldsymbol{M}_{k}=\boldsymbol{B}_{k}\approx\boldsymbol{H}_{k}$ such that still a superlinear convergence rate can be achieved
$\boldsymbol{M}(\boldsymbol{\theta})=\boldsymbol{M}_{1}(\boldsymbol{\theta})+\boldsymbol{M}_{2}(\boldsymbol{\theta})$ (20)
where the matrices $\boldsymbol{M}_{1}(\boldsymbol{\theta}),\boldsymbol{M}_{2}(\boldsymbol{\theta})\in\mathbb{R}^{n_{\boldsymbol{\theta}}\times n_{\boldsymbol{\theta}}}$ are defined as
$\boldsymbol{M}_{1}(\boldsymbol{\theta})$ $=\mathbb{E}_{\boldsymbol{s}}\left[\nabla_{\boldsymbol{\theta}}^{2}\pi_{\boldsymbol{\theta}}(\boldsymbol{s})\otimes\nabla_{\boldsymbol{a}}Q^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s},\boldsymbol{a})|_{\boldsymbol{a}=\pi_{\boldsymbol{\theta}}(\boldsymbol{s})}\right]$ (21a)
$\boldsymbol{M}_{2}(\boldsymbol{\theta})$ $=\mathbb{E}_{\boldsymbol{s}}\left[\nabla_{\boldsymbol{\theta}}\pi_{\boldsymbol{\theta}}^{\top}(\boldsymbol{s})\nabla_{\boldsymbol{a}}^{2}Q^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s},\boldsymbol{a})|_{\boldsymbol{a}=\pi_{\boldsymbol{\theta}}(\boldsymbol{s})}\nabla_{\boldsymbol{\theta}}\pi_{\boldsymbol{\theta}}(\boldsymbol{s})\right]$ (21b)
with $\nabla_{\boldsymbol{\theta}}^{2}\pi_{\boldsymbol{\theta}}(\boldsymbol{s})\in\mathbb{R}^{n_{\boldsymbol{a}}\times n_{\boldsymbol{\theta}}\times n_{\boldsymbol{\theta}}}$ being the second-order Jacobian tensor of the policy with respect to the policy parameters, $\nabla_{\boldsymbol{a}}^{2}Q^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s},\boldsymbol{a})\in\mathbb{R}^{n_{\boldsymbol{a}}\times n_{\boldsymbol{a}}}$ denoting the Hessian of the action-value function with respect to the taken action, and the operator $\otimes:\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}\times\mathbb{R}^{n_{1}}\rightarrow\mathbb{R}^{n_{2}\times n_{3}}$ denoting the tensor-vector product, which we define as
$\boldsymbol{\mathrm{T}}\otimes\boldsymbol{v}=\sum_{i=1}^{n_{1}}\boldsymbol{\mathrm{T}}_{[i,:,:]}\boldsymbol{v}_{i}$ (22)
with $\boldsymbol{\mathrm{T}}_{[i,:,:]}\in\mathbb{R}^{n_{2}\times n_{3}}$ denoting the $i$-th matrix slice of the general tensor $\boldsymbol{\mathrm{T}}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ and $\boldsymbol{v}_{i}\in\mathbb{R}$ denoting the $i$-th element of the general vector $\boldsymbol{v}\in\mathbb{R}^{n_{1}}$.
In the context of RL, the policy Hessian approximation (20) is computed using an expected value like the policy gradient. This expected value must be estimated empirically with a finite amount of samples similar to (13), resulting in $\tilde{\boldsymbol{B}}_{k}\approx\boldsymbol{B}_{k}$. Subsampled Newton methods form the subcategory of Newton methods that consider optimization with noisy gradient $\tilde{\boldsymbol{g}}_{k}$ and Hessian estimates $\tilde{\boldsymbol{B}}_{k}$. Detailed reviews on subsampled Newton methods are provided in Bollapragada et al. (2019) and Bottou et al. (2018). Analogous to (14) and (19), the update rule for subsampled Newton methods is derived as
$\boldsymbol{p}_{k}=-\tilde{\boldsymbol{B}}_{k}^{-1}\tilde{\boldsymbol{g}}_{k}.$ (23)
One central idea in subsampled Newton methods is that the quality of the obtained update in (23) is mainly determined by the accuracy of the gradient estimate $\tilde{\boldsymbol{g}}_{k}$. The accuracy of the Hessian estimate only has a minor influence on the proposed update rule as long as it is of sufficient quality. In many applications the Hessian estimate can be improved by taking large enough samples sizes. In practical RL approaches however, the Hessian estimation is more challenging because the action-value function $Q^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s},\boldsymbol{a})$ is typically approximated. Insufficient action-value function approximation can affect the Hessian estimate severely leading to unstable updates. In the context of general large-scale optimization, approaches like AdaHessian (Yao et al., 2021) or Sophia (Liu et al., 2024) aim to alleviate instabilities due to insufficient Hessian estimation with the use of momentum on the Hessian estimate. Both approaches constrain themselves to vectorial Hessian approximations as they are particularly designed for large scale optimization and do not consider a good initialization of the exponentially averaged Hessian estimate.
2.3 Parameterized model predictive control as policy in reinforcement learning
RL agents act according to their policy. Most works in the area of RL focus on approximating the optimal policy using linear functions or neural networks (NNs). Despite having a good interpretability, linear functions are often limited in their expressiveness. NNs on the other side can represent almost any desired policy but lack interpretability due to their black-box mapping. In addition to that, if the NN cannot be pretrained, the NN parameters are usually initialized randomly, leading to an excessive amount of required training data. In this work, we focus on the approximation of the optimal policy via a parameterized MPC scheme.
MPC is a model-based optimal control approach that computes a sequence of optimal control actions $\boldsymbol{\mathrm{u}}^{\star}=({\boldsymbol{u}_{0}^{\star}},\ldots,{\boldsymbol{u}_{N-1}^{\star}})$ with $\boldsymbol{u}_{i}\in\mathcal{A}$ along a prediction horizon $N\in\mathbb{N}$ based on an internal system model such that a control objective is optimized and process constraint are satisfied. Gros and Zanon (2020) show that the following parameterized MPC optimization problem can be used as a policy in RL to capture the optimal policy
$\boldsymbol{\mathrm{u}}^{\star}(\boldsymbol{s})=\arg\min_{\boldsymbol{\mathrm{u}}}\leavevmode\nobreak\ \leavevmode\nobreak$ $V_{\mathrm{f},\boldsymbol{\theta}}(\boldsymbol{x}_{N})+\boldsymbol{w}_{\mathrm{f}}^{\top}\boldsymbol{\sigma}_{N}$
$+\sum_{i=0}^{N-1}\ell_{\boldsymbol{\theta}}(\boldsymbol{x}_{i},\boldsymbol{u}_{i})+\boldsymbol{w}^{\top}\boldsymbol{\sigma}_{i}$ (24a)
$\mathrm{s.t.}\leavevmode\nobreak\ \leavevmode\nobreak$ $\boldsymbol{x}_{i+1}=f^{\mathrm{MPC}}_{\boldsymbol{\theta}}(\boldsymbol{x}_{i},\boldsymbol{u}_{i}),\leavevmode\nobreak\ \boldsymbol{x}_{0}=\boldsymbol{s}$ (24b)
$\boldsymbol{u}_{\mathrm{lb}}\leq\boldsymbol{u}_{i}\leq\boldsymbol{u}_{\mathrm{ub}}$ (24c)
$g_{\boldsymbol{\theta}}(\boldsymbol{x}_{i},\boldsymbol{u}_{i})\leq\boldsymbol{\sigma}_{i},\quad\boldsymbol{\sigma}_{i}\geq 0$ (24d)
$h_{\boldsymbol{\theta}}(\boldsymbol{x}_{i},\boldsymbol{u}_{i})=0.$ (24e)
In (24), the objective function (24a) is minimized. It consists of the sum of stage costs $\ell_{\boldsymbol{\theta}}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ starting from the initial condition at $i=0$ up to the second to last point of the prediction horizon at $i=N-1$, the terminal cost at the end of the prediction $V_{\mathrm{f},\boldsymbol{\theta}}:\mathcal{S}\rightarrow\mathbb{R}$, and penalties for soft constraints. The system dynamics are accounted for via the system equations $f^{\mathrm{MPC}}_{\boldsymbol{\theta}}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ starting from the initial state $\boldsymbol{x}_{0}$ as stated in (24b). The predicted trajectory does not need to be the same as in the MDP since $f^{\mathrm{MPC}}_{\boldsymbol{\theta}}$ is usually only an approximation of the true system dynamics (1). Maximum and minimum values for the control action are posed via (24c). Process equality and inequality constraints are ensured via $h_{\boldsymbol{\theta}}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{n_{h}}$ and $g_{\boldsymbol{\theta}}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{n_{g}}$ as shown in (24e) and (24d) respectively. To always ensure feasibility, the inequality constraints need to be relaxed to soft constraints. All functions with the index $\boldsymbol{\theta}$ are parameterizable and can be adapted by the RL algorithm. Note that to distinguish the states and actions within the optimal control problem (24) from those within the MDP, we introduce the states and actions within the MPC as $\boldsymbol{x}$ and $\boldsymbol{u}$. After (24) is solved, the MPC policy is derived by taking the first element of the sequence of optimal control actions $\boldsymbol{\mathrm{u}}^{\star}$, so $\pi_{\boldsymbol{\theta}}(\boldsymbol{s})=\boldsymbol{u}_{0}^{\star}$.
2.3.1 Parametric NLP sensitivities
In order to update the policy parameters with a policy optimization method, the Jacobian of the policy with respect to its parameters needs to be computed. When using MPC as a policy, the derivative of the policy equals the derivative of the solution of the optimization problem, which is only computable explicitly in rare cases. The MPC optimization problem (24) can be categorized as a parametric constrained nonlinear program (NLP). The derivative of the optimal solution with respect to the parameters of the NLP is also often referred to as the (parametric) NLP sensitivities. We will show next, how the NLP sensitivities can be computed.
We first introduce a general constrained NLP of the form
$\boldsymbol{z}^{\star}(\boldsymbol{\zeta})=\arg\min_{\boldsymbol{z}}\leavevmode\nobreak\ \leavevmode\nobreak$ $\Phi(\boldsymbol{z};\boldsymbol{\zeta})$ (25a)
$\mathrm{s.t.}\leavevmode\nobreak\ \leavevmode\nobreak$ $g(\boldsymbol{z};\boldsymbol{\zeta})\leq 0$ (25b)
$h(\boldsymbol{z};\boldsymbol{\zeta})=0$ (25c)
where $\boldsymbol{z}\in\mathbb{R}^{n_{\boldsymbol{z}}}$ denotes the decision variables, $\boldsymbol{\zeta}\in\mathbb{R}^{n_{\boldsymbol{\zeta}}}$ denotes the changeable parameters, $\Phi:\mathbb{R}^{n_{\boldsymbol{z}}}\times\mathbb{R}^{n_{\boldsymbol{\zeta}}}\rightarrow\mathbb{R}$ denotes the scalar objective function, $g:\mathbb{R}^{n_{\boldsymbol{z}}}\times\mathbb{R}^{n_{\boldsymbol{\zeta}}}\rightarrow\mathbb{R}^{n_{g}}$ denotes the inequality constraints, and $h:\mathbb{R}^{n_{\boldsymbol{z}}}\times\mathbb{R}^{n_{\boldsymbol{\zeta}}}\rightarrow\mathbb{R}^{n_{h}}$ denotes the equality constraints. The optimal solution $\boldsymbol{z}^{\star}(\boldsymbol{\zeta})$ of (25) changes with the optimization parameters $\boldsymbol{\zeta}$. Except for rare cases, this functional relationship is typically unknown. We introduce the Lagrangian $\mathcal{L}:\mathbb{R}^{n_{\boldsymbol{z}}}\times\mathbb{R}^{n_{g}}\times\mathbb{R}^{n_{h}}\times\mathbb{R}^{n_{\boldsymbol{\zeta}}}\rightarrow\mathbb{R}$ of (25) as
$\mathcal{L}(\boldsymbol{z},\boldsymbol{\lambda},\boldsymbol{\nu};\boldsymbol{\zeta})=\Phi(\boldsymbol{z};\boldsymbol{\zeta})+\boldsymbol{\lambda}^{\top}g(\boldsymbol{z};\boldsymbol{\zeta})+\boldsymbol{\nu}^{\top}h(\boldsymbol{z};\boldsymbol{\zeta})$ (26)
where $\boldsymbol{\lambda}\in\mathbb{R}^{n_{g}}$ and $\boldsymbol{\nu}\in\mathbb{R}^{n_{h}}$ are the Lagrange multipliers, also known as dual variables, associated to the inequality and equality constraints respectively. To simplify notation, we introduce the stacked primal-dual vector $\boldsymbol{\chi}^{\top}=(\boldsymbol{z}^{\top},\boldsymbol{\lambda}^{\top},\boldsymbol{\nu}^{\top})$ with $\boldsymbol{\chi}\in\mathbb{R}^{n_{\boldsymbol{\chi}}}$ and $n_{\boldsymbol{\chi}}=n_{\boldsymbol{z}}+n_{g}+n_{h}$. Let $\boldsymbol{\chi}^{\star}(\boldsymbol{\zeta})$ denote an optimal primal-dual vector solving (25) and satisfying the Karush-Kuhn-Tucker (KKT) conditions (Nocedal and Wright, 2006), and let $g$ satisfy the linear-independency constraint qualifications (Nocedal and Wright, 2006), then the equalities of the KKT conditions form an implicit function $F:\mathbb{R}^{n_{\boldsymbol{\chi}}}\times\mathbb{R}^{n_{\boldsymbol{\zeta}}}\rightarrow\mathbb{R}^{n_{\boldsymbol{\chi}}}$ of the form
$F(\boldsymbol{\chi}^{\star}(\boldsymbol{\zeta});\boldsymbol{\zeta})=\begin{pmatrix}\nabla_{\boldsymbol{z}}\mathcal{L}(\boldsymbol{\chi};\boldsymbol{\zeta})|_{\boldsymbol{\chi}=\boldsymbol{\chi}^{\star}(\boldsymbol{\zeta})}\\ h(\boldsymbol{z}^{\star}(\boldsymbol{\zeta});\boldsymbol{\zeta})\\ \boldsymbol{\lambda}^{\star}(\boldsymbol{\zeta})\odot g(\boldsymbol{z}^{\star}(\boldsymbol{\zeta});\boldsymbol{\zeta})\end{pmatrix}=0$ (27)
where the operator $\odot$ denotes the Hadamard (elementwise) product. In order to obtain the first-order NLP sensitivities of the optimal primal-dual solution with respect to the NLP parameters $\nabla_{\boldsymbol{\zeta}}\boldsymbol{\chi}^{\star}(\boldsymbol{\zeta})\in\mathbb{R}^{n_{\boldsymbol{\chi}}\times n_{\boldsymbol{\zeta}}}$, (27) can be implicitly differentiated almost everywhere (Fiacco and Ishizuka, 1990) forming the following linear system
$\nabla_{\boldsymbol{\chi}^{\star}}F(\boldsymbol{\chi}^{\star};\boldsymbol{\zeta})\,\nabla_{\boldsymbol{\zeta}}\boldsymbol{\chi}^{\star}(\boldsymbol{\zeta})=-\nabla_{\boldsymbol{\zeta}}F(\boldsymbol{\chi}^{\star};\boldsymbol{\zeta})$ (28)
with the coefficient matrix $\nabla_{\boldsymbol{\chi}^{\star}}F(\boldsymbol{\chi}^{\star};\boldsymbol{\zeta})\in\mathbb{R}^{n_{\boldsymbol{\chi}}\times n_{\boldsymbol{\chi}}}$ and the right-hand-side matrix $-\nabla_{\boldsymbol{\zeta}}F(\boldsymbol{\chi}^{\star};\boldsymbol{\zeta})\in\mathbb{R}^{n_{\boldsymbol{\chi}}\times n_{\boldsymbol{\zeta}}}$. The first-order NLP sensitivities can therefore be obtained by solving (28) for $\nabla_{\boldsymbol{\zeta}}\boldsymbol{\chi}^{\star}(\boldsymbol{\zeta})$. In the context of MPC as policy in RL, Brandner and Lucia (2024) show that the second-order NLP sensitivities can be obtained in a similar fashion and also boil down to the solution of a linear system. However the right-hand-side matrix requires third partial derivatives of the Lagrangian with respect to full primal-dual vector. This can render the computation intractable, for instance if the number of primal-dual variables $\boldsymbol{\chi}$ is large, or if highly nonlinear terms appear in the NLP such that the third partial derivatives become extremely convoluted.
To compute the deterministic policy gradient (12) and the respective terms of the deterministic policy Hessian (20), the NLP (25) must be cast into (24) by setting $\boldsymbol{z}^{\star}=\boldsymbol{\mathrm{u}}^{\star}$ and $\boldsymbol{\zeta}=\boldsymbol{\theta}$. The relevant sensitivities $\nabla_{\boldsymbol{\theta}}\pi_{\boldsymbol{\theta}}(\boldsymbol{s})=\nabla_{\boldsymbol{\theta}}\boldsymbol{u}_{0,\boldsymbol{\theta}}^{\star}(\boldsymbol{s})$ must then be extracted from the full sensitivity matrix.
3 Robust computationally efficient Gauss-Newton policy optimization
Using MPC as policy within RL can significantly alleviate some of the main challenges of RL based on NN policies: it provides a good initial policy and can deal with hard constraints in real-time by solving an optimization problem. Although the sparse parameterization of MPC policies enables the use of second-order optimization methods, the training still remains computationally challenging, as Hessian estimates such as (20) rely on the second-order NLP sensitivities. For that reason, we propose in Section 3.1 a novel Gauss-Newton approximation of the deterministic policy Hessian that retains superlinear convergence while avoiding the computation of the second-order NLP sensitivities. We then present in Section 3.2 a new approach to introduce momentum in the Gauss-Newton update to increase numerical stability.
3.1 Gauss-Newton approximation of the deterministic policy Hessian
In order to avoid the computation of the second-order NLP sensitivities, we will derive an approximation of the deterministic policy Hessian that does not depend on those, but still gives superlinear convergence guarantees. Lets first consider the behavior of the action-value function of the optimal policy around the the optimal action $\boldsymbol{a}=\pi^{\star}(\boldsymbol{s})$.
Lemma 1
Let the action-value function of the optimal policy $Q^{\pi^{\star}}(\boldsymbol{s},\boldsymbol{a})$ be differentiable in $\boldsymbol{a}$ at $\boldsymbol{a}=\pi_{\boldsymbol{\theta}^{\star}}(\boldsymbol{s})$ for all $\boldsymbol{s}\in\mathcal{S}$, then the gradient evaluated at the optimal action is zero such that
$\nabla_{\boldsymbol{a}}Q^{\pi^{\star}}(\boldsymbol{s},\boldsymbol{a})|_{\boldsymbol{a}=\pi^{\star}(\boldsymbol{s})}=0,\quad\forall\boldsymbol{s}\in\mathcal{S}.$ (29)
{pf}
Due to its definition, the optimal action $\boldsymbol{a}=\pi^{\star}(\boldsymbol{s})$ has a better or at least equal value assigned to it when following the optimal policy $\pi^{\star}$ than any other action, as it is defined in (8). Therefore, $Q^{\pi^{\star}}(\boldsymbol{s},\pi^{\star}(\boldsymbol{s}))$ is a maximum. Since $Q^{\pi^{\star}}(\boldsymbol{s},\boldsymbol{a})$ is assumed to be continuously differentiable at $\boldsymbol{a}=\pi^{\star}(\boldsymbol{s})$, the first-order optimality condition is given as $\nabla_{\boldsymbol{a}}Q^{\pi^{\star}}(\boldsymbol{s},\boldsymbol{a})|_{\boldsymbol{a}=\pi^{\star}(\boldsymbol{s})}=0$. $\qed$
Corollary 1
Let $Q^{\pi^{\star}}(\boldsymbol{s},\boldsymbol{a})$ be twice differentiable in $\boldsymbol{a}$ at $\boldsymbol{a}=\pi^{\star}(\boldsymbol{s})$ then the Hessian with respect to $\boldsymbol{a}$ at $\boldsymbol{a}=\pi^{\star}(\boldsymbol{s})$ is negative semi-definite
$\nabla_{\boldsymbol{a}}^{2}Q^{\pi^{\star}}(\boldsymbol{s},\boldsymbol{a})|_{\boldsymbol{a}=\pi^{\star}(\boldsymbol{s})}\preceq 0.$ (30)
{pf}
To be a unique maximum, $\nabla_{\boldsymbol{a}}^{2}Q^{\pi^{\star}}(\boldsymbol{s},\boldsymbol{a})|_{\boldsymbol{a}=\pi^{\star}(\boldsymbol{s})}$ must be strictly negative definite. Since $Q^{\pi^{\star}}(\boldsymbol{s},\boldsymbol{a})$ can also be locally constant for $\boldsymbol{a}$ in the neighborhood of $\pi^{\star}(\boldsymbol{s})$ if the optimal policy is not unique, the curvature can also be zero. As the result $\nabla_{\boldsymbol{a}}^{2}Q^{\pi^{\star}}(\boldsymbol{s},\boldsymbol{a})|_{\boldsymbol{a}=\pi^{\star}(\boldsymbol{s})}$ is negative semi-definite. $\qed$
Kordabad et al. (2022) proves that the approximation of the deterministic policy Hessian stated in (20) converges to the true deterministic policy Hessian for $\boldsymbol{\theta}\rightarrow\boldsymbol{\theta}^{\star}$. We show that even when neglecting the first part $\boldsymbol{M}_{1}(\boldsymbol{\theta})$ of the approximation in (20), $\boldsymbol{M}_{2}(\boldsymbol{\theta})\approx\nabla_{\boldsymbol{\theta}}^{2}J(\boldsymbol{\theta})$ converges to the true deterministic policy Hessian for $\boldsymbol{\theta}\rightarrow\boldsymbol{\theta}^{\star}$ with the benefit that the second-order NLP sensitivities can be omitted. A Hessian approximation with the structure of $\boldsymbol{M}_{2}(\theta)$ is also referred to as a Gauss-Newton approximation (Bottou et al., 2018). We first state the following assumption.
Assumption 1
The parameterized policy $\pi_{\boldsymbol{\theta}}(\boldsymbol{s})$ is rich enough so there exists $\boldsymbol{\theta}^{\star}$ such that $\pi_{\boldsymbol{\theta}^{\star}}(\boldsymbol{s})=\pi^{\star}(\boldsymbol{s})$.
Theorem 1
Let Assumption 1 hold and further let $\nabla_{\boldsymbol{\theta}}^{2}\pi_{\boldsymbol{\theta}}(\boldsymbol{s})$ and $\nabla_{\boldsymbol{a}}Q^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s},\boldsymbol{a})|_{\boldsymbol{a}=\pi_{\boldsymbol{\theta}}(\boldsymbol{s})}$ be bounded for all $\boldsymbol{\theta}$, then the approximation $\boldsymbol{M}_{2}(\boldsymbol{\theta})\approx\nabla_{\boldsymbol{\theta}}^{2}J(\boldsymbol{\theta})$ converges for $\boldsymbol{\theta}\rightarrow\boldsymbol{\theta}^{\star}$ to the true deterministic policy Hessian at the optimal policy, so $\boldsymbol{M}_{2}(\boldsymbol{\theta}^{\star})=\nabla_{\boldsymbol{\theta}}^{2}J(\boldsymbol{\theta})|_{\boldsymbol{\theta}=\boldsymbol{\theta}^{\star}}$ and moreover
$\lim_{\boldsymbol{\theta}\rightarrow\boldsymbol{\theta}^{\star}}\boldsymbol{M}_{1}(\boldsymbol{\theta})=0.$ (31)
{pf}
We define the auxiliary function $\varphi_{\boldsymbol{\theta}}:\mathcal{S}\rightarrow\mathbb{R}^{n_{\boldsymbol{\theta}}\times n_{\boldsymbol{\theta}}}$
$\varphi_{\boldsymbol{\theta}}(\boldsymbol{s})=\nabla_{\boldsymbol{\theta}}^{2}\pi_{\boldsymbol{\theta}}(\boldsymbol{s})\otimes\nabla_{\boldsymbol{a}}Q^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s},\boldsymbol{a})|_{\boldsymbol{a}=\pi_{\boldsymbol{\theta}}(\boldsymbol{s})}.$ (32)
Since $\nabla_{\boldsymbol{\theta}}^{2}\pi_{\boldsymbol{\theta}}(\boldsymbol{s})$ and $\nabla_{\boldsymbol{a}}Q^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s},\boldsymbol{a})|_{\boldsymbol{a}=\pi_{\boldsymbol{\theta}}(\boldsymbol{s})}$ are bounded by assumption, $\varphi_{\boldsymbol{\theta}}(\boldsymbol{s})$ is also bounded and can even be bounded uniformly. We compute the limit of $\varphi_{\boldsymbol{\theta}}(\boldsymbol{s})$ for $\boldsymbol{\theta}\rightarrow\boldsymbol{\theta}^{\star}$ to be
$\lim_{\boldsymbol{\theta}\rightarrow\boldsymbol{\theta}^{\star}}\varphi_{\boldsymbol{\theta}}(\boldsymbol{s})$
$=$ $\lim_{\boldsymbol{\theta}\rightarrow\boldsymbol{\theta}^{\star}}\nabla_{\boldsymbol{\theta}}^{2}\pi_{\boldsymbol{\theta}}(\boldsymbol{s})\otimes\nabla_{\boldsymbol{a}}Q^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s},\boldsymbol{a})|_{\boldsymbol{a}=\pi_{\boldsymbol{\theta}}(\boldsymbol{s})}$
$=$ $\lim_{\boldsymbol{\theta}\rightarrow\boldsymbol{\theta}^{\star}}\nabla_{\boldsymbol{\theta}}^{2}\pi_{\boldsymbol{\theta}}(\boldsymbol{s})\otimes\lim_{\boldsymbol{\theta}\rightarrow\boldsymbol{\theta}^{\star}}\nabla_{\boldsymbol{a}}Q^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s},\boldsymbol{a})|_{\boldsymbol{a}=\pi_{\boldsymbol{\theta}}(\boldsymbol{s})}$
$=$ $\nabla_{\boldsymbol{\theta}}^{2}\pi_{\boldsymbol{\theta}^{\star}}(\boldsymbol{s})\otimes 0=0,$ (33)
as it follows from Theorem 1 that $\nabla_{\boldsymbol{a}}Q^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s},\boldsymbol{a})|_{\boldsymbol{a}=\pi_{\boldsymbol{\theta}}(\boldsymbol{s})}$ vanishes at $\boldsymbol{\theta}^{\star}$ and $\nabla_{\boldsymbol{\theta}}^{2}\pi_{\boldsymbol{\theta}}(\boldsymbol{s})$ is assumed to be bounded.
From definition (21a), it follows that
$\lim_{\boldsymbol{\theta}\rightarrow\boldsymbol{\theta}^{\star}}\boldsymbol{M}_{1}(\boldsymbol{\theta})$ $=\lim_{\boldsymbol{\theta}\rightarrow\boldsymbol{\theta}^{\star}}\mathbb{E}_{\boldsymbol{s}}\left[\varphi_{\boldsymbol{\theta}}(\boldsymbol{s})\right]$
$=\lim_{\boldsymbol{\theta}\rightarrow\boldsymbol{\theta}^{\star}}\int_{\mathcal{S}}\rho_{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s})\varphi_{\boldsymbol{\theta}}(\boldsymbol{s})\,\mathrm{d}\boldsymbol{s}.$ (34)
Here, $\rho_{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s}):\mathcal{S}\rightarrow[0,\infty)$ denotes the probability density that state $\boldsymbol{s}$ is encountered under closed-loop operation of $\pi_{\boldsymbol{\theta}}$. Since $\rho_{\pi_{\boldsymbol{\theta}}}$ and $\varphi_{\boldsymbol{\theta}}(\boldsymbol{s})$ can be uniformly bounded, the dominated convergence theorem (Bronštejn et al., 2016) can be applied so the limit operator can be moved inside the integral. It follows that
$\lim_{\boldsymbol{\theta}\rightarrow\boldsymbol{\theta}^{\star}}\int_{\mathcal{S}}\rho_{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s})\varphi_{\boldsymbol{\theta}}(\boldsymbol{s})\,\mathrm{d}\boldsymbol{s}$
$=$ $\int_{\mathcal{S}}\lim_{\boldsymbol{\theta}\rightarrow\boldsymbol{\theta}^{\star}}\rho_{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s})\varphi_{\boldsymbol{\theta}}(\boldsymbol{s})\,\mathrm{d}\boldsymbol{s}$
$=$ $\int_{\mathcal{S}}\lim_{\boldsymbol{\theta}\rightarrow\boldsymbol{\theta}^{\star}}\rho_{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s})\cdot\lim_{\boldsymbol{\theta}\rightarrow\boldsymbol{\theta}^{\star}}\varphi_{\boldsymbol{\theta}}(\boldsymbol{s})\,\mathrm{d}\boldsymbol{s}$
$=$ $\int_{\mathcal{S}}\rho_{\pi_{\boldsymbol{\theta}^{\star}}}(\boldsymbol{s})\cdot 0\,\mathrm{d}\boldsymbol{s}=\int_{\mathcal{S}}0\,\mathrm{d}\boldsymbol{s}=0,$ (35)
where the equality in the last line holds because of (3.1), proving that $\boldsymbol{M}_{1}(\boldsymbol{\theta})$ vanishes if $\rho_{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s})$ is bounded.
Corollary 2
Let Assumption 1 hold and let $\boldsymbol{M}_{2}(\boldsymbol{\theta})$ be defined for $\boldsymbol{\theta}$ in the neighborhood of $\boldsymbol{\theta}^{\star}$, then $\boldsymbol{M}_{2}(\boldsymbol{\theta})$ is negative semi-definite in the neighborhood of $\boldsymbol{\theta}^{\star}$
$\boldsymbol{M}_{2}(\boldsymbol{\theta})\preceq 0.$ (36)
{pf}
Since $\nabla_{\boldsymbol{a}}^{2}Q^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s},\boldsymbol{a})|_{\boldsymbol{a}=\pi_{\boldsymbol{\theta}}(\boldsymbol{s})}$ is negative semi-definite for $\boldsymbol{\theta}$ in the neighborhood of $\boldsymbol{\theta}^{\star}$, it follows from the fact that the result of a left and right multiplication with any matrix and its transposed to a negative semi-definite matrix is again a negative semi-definite matrix (Bronštejn et al., 2016) so
$\nabla_{\boldsymbol{\theta}}\pi_{\boldsymbol{\theta}}^{\top}(\boldsymbol{s})\nabla_{\boldsymbol{a}}^{2}Q^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s},\boldsymbol{a})|_{\boldsymbol{a}=\pi_{\boldsymbol{\theta}}(\boldsymbol{s})}\nabla_{\boldsymbol{\theta}}\pi_{\boldsymbol{\theta}}(\boldsymbol{s})\preceq 0.$ (37)
As the expected value in the definition of $\boldsymbol{M}_{2}(\boldsymbol{\theta})$ boils down to an integral over negative semi-definite matrices, the final result is also negative semi-definite. $\qed$
Finally, we proof that choosing the approximate Hessian $\boldsymbol{B}_{k}$ as $\boldsymbol{M}_{2,k}=\boldsymbol{B}_{k}$ still guarantees superlinear convergence. For this, consider the condition for quasi-Newton methods for superlinear convergence.
Theorem 2
Suppose that $J:\mathbb{R}^{n_{\boldsymbol{\theta}}}\rightarrow\mathbb{R}$ is twice differentiable. Consider the iteration $\boldsymbol{\theta}_{k+1}=\boldsymbol{\theta}_{k}+\boldsymbol{p}_{k}$ with $\boldsymbol{p}_{k}$ defined as in (19) and let the sequence of iterations $\{\boldsymbol{\theta}_{k}\}$ converge to $\boldsymbol{\theta}^{\star}$ such that $\nabla_{\boldsymbol{\theta}}J(\boldsymbol{\theta})|_{\boldsymbol{\theta}=\boldsymbol{\theta}^{\star}}=0$ and $\nabla_{\boldsymbol{\theta}}^{2}J(\boldsymbol{\theta})|_{\boldsymbol{\theta}=\boldsymbol{\theta}^{\star}}$ is negative definite. Then $\{\boldsymbol{\theta}_{k}\}$ converges superlinearly if
$\lim_{k\rightarrow\infty}\frac{\|\left(\boldsymbol{B}_{k}-\nabla_{\boldsymbol{\theta}}^{2}J(\boldsymbol{\theta})|_{\boldsymbol{\theta}=\boldsymbol{\theta}^{\star}}\right)\boldsymbol{p}_{k}\|}{\|\boldsymbol{p}_{k}\|}=0.$ (38)
{pf}
See Theorem 3.7. in Nocedal and Wright (2006).
Corollary 3
Let Assumption 1 hold, lets further assume that all assumptions from Theorem 2 hold, and assume that the optimal policy is unique, then the sequence of iterations $\{\boldsymbol{\theta}_{k}\}$ in order to solve (9) converges superlinearly to $\boldsymbol{\theta}^{\star}$ if $\boldsymbol{B}_{k}=\boldsymbol{M}_{2,k}$ and update rule (19) is used.
{pf}
Since a unique optimal policy is assumed, $\boldsymbol{M}_{2}(\theta^{\star})=\nabla_{\boldsymbol{\theta}}^{2}J(\boldsymbol{\theta})|_{\boldsymbol{\theta}=\boldsymbol{\theta}^{\star}}$ is strictly negative definite. Further $\boldsymbol{B}_{k}=\boldsymbol{M}_{2,k}$ converges to $\nabla_{\boldsymbol{\theta}}^{2}J(\boldsymbol{\theta})_{\boldsymbol{\theta}=\boldsymbol{\theta}^{\star}}$ as shown in Theorem 1. Therefore condition (38) is satisfied. $\qed$
3.2 Robust Gauss-Newton policy optimization with momentum
If the action-value function $Q^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s},\boldsymbol{a})$ of the current policy $\pi_{\boldsymbol{\theta}}(\boldsymbol{s})$ is explicitly known, and the expected values could be computed perfectly, the proposed approximate Newton method would converge superlinearly. However, in reality, the action-value function of the current policy must be approximated. Further, the expected value must be computed empirically by drawing a finite amount of samples. Both can lead to noisy estimates of the policy gradient $\tilde{\boldsymbol{g}}_{k}\approx\boldsymbol{g}_{k}$ and the approximate policy Hessian $\tilde{\boldsymbol{B}}_{k}\approx\boldsymbol{B}_{k}$ preventing accurate parameter updates and causing potential instability. A common choice to mitigate the effect of noisy estimates within first-order stochastic optimization is the use of momentum-based approaches as introduced in Section 2.2.1. While second-order updates are usually robust with respect to poorly estimated Hessian approximations (Bollapragada et al., 2019; Bottou et al., 2018), the estimates still need to be of reasonable accuracy. A single poor Hessian estimation can already cause instability, for example if the eigenvalues that are estimated are orders of magnitude wrong, causing an update that deteriorates the parameter iterates to values that are far from optimal. We show next how the ideas of momentum can be transferred to second-order approaches by applying some adaptations.
Analogous to the update rule of stochastic gradient ascent with momentum (17), we introduce the update rule
$\boldsymbol{p}_{k}=-\alpha\hat{\boldsymbol{D}}_{k}^{-1}\hat{\boldsymbol{m}}_{k}$ (39)
in alignment with the subsampled Newton update rule (23). Here, $\hat{\boldsymbol{m}}_{k}$ denotes the unbiased averaged gradient estimate as described by the momentum-based updates in (15) and (16). The unbiased averaged Hessian estimate that we derive in this work is denoted by $\hat{\boldsymbol{D}}_{k}\in\mathbb{R}^{n_{\boldsymbol{\theta}}\times n_{\boldsymbol{\theta}}}$. Lets define the biased averaged Hessian estimate $\boldsymbol{D}_{k}\in\mathbb{R}^{n_{\boldsymbol{\theta}}\times n_{\boldsymbol{\theta}}}$ recursively as
$\boldsymbol{D}_{k}=\eta\boldsymbol{D}_{k-1}+(1-\eta)\tilde{\boldsymbol{B}}_{k}$ (40)
with $0\leq\eta<1$ denoting the exponential averaging factor for Hessian estimation, $\boldsymbol{D}_{k-1}$ denoting the biased averaged Hessian estimation of the previous iteration, and $\tilde{\boldsymbol{B}}_{k}$ being the local Hessian estimate at the current iteration. Note that we do not specify $\tilde{\boldsymbol{B}}_{k}$ as it can be any Hessian approximation such as $\tilde{\boldsymbol{B}}_{k}=\tilde{\boldsymbol{M}}_{k}$ or $\tilde{\boldsymbol{B}}_{k}=\tilde{\boldsymbol{M}}_{2,k}$.
As in the case of the gradient momentum, the matrix $\boldsymbol{D}_{0}$ must be initialized. Although it may be appealing to initialize with a zero matrix $\boldsymbol{D}_{0}=0$, it turns out to be impractical because especially in the first few steps of the optimization, $\boldsymbol{D}_{k}$ would be severely influenced by inaccurate local estimates $\tilde{\boldsymbol{B}}_{k}$. Consider for instance the scenario that the eigenvalues $\lambda(\tilde{\boldsymbol{B}}_{1})$ of $\tilde{\boldsymbol{B}}_{1}$ are in the order of magnitude of $\lambda_{i}(\tilde{\boldsymbol{B}}_{1})\sim 1$ whereas the eigenvalues of $\boldsymbol{B}_{1}$ are rather in the order of magnitude of $\lambda_{i}(\boldsymbol{B}_{1})\sim 10^{3}$, the inverse of the Hessian in (39) would scale the gradient roughly by the inverse of the eigenvalues, which is $1/\lambda_{i}(\tilde{\boldsymbol{B}}_{1})\sim 1$ although it should rather be $1/\lambda_{i}(\boldsymbol{B}_{1})\sim 10^{-3}$. To avoid such undesirable behavior, we propose to initialize with a multiple of the identity matrix
$\boldsymbol{D}_{0}=-\omega^{-1}\boldsymbol{I}$ (41)
where $\omega>0$ can be understood as an initial learning rate since $\boldsymbol{D}_{0}^{-1}=-\omega\boldsymbol{I}$ would scale each element of the gradient by $\omega$ when applied in (39). Further, $\omega^{-1}$ should be chosen as a rough estimate of the largest absolute eigenvalue of $\boldsymbol{B}_{1}$.
To account for the initialization bias towards $\boldsymbol{D}_{0}$, we derive a scaling function similar to (16) that allows fast adjustments to newly obtained local Hessian estimates $\tilde{\boldsymbol{B}}_{k}$ and still maintains the added robustness from $\boldsymbol{D}_{0}$. Writing the recursive definition (40) explicitly and using $-\omega^{-1}\boldsymbol{I}=(1-\eta)\tilde{\boldsymbol{B}}_{0}$ gives
$\boldsymbol{D}_{k}=$ $-\eta^{k}\omega^{-1}\boldsymbol{I}+(1-\eta)\sum_{i=1}^{k}\eta^{k-i}\tilde{\boldsymbol{B}_{i}}$
$=$ $(1-\eta)\sum_{i=0}^{k}\eta^{k-i}\tilde{\boldsymbol{B}_{i}}.$ (42)
This expression aligns closely to results from Kingma and Ba (2015) with the exception that the sum starts with $i=0$ instead of $i=1$. Following a similar derivation as in Kingma and Ba (2015), we derive the following corrected exponential moving average $\hat{\boldsymbol{D}}_{k}$
$\hat{\boldsymbol{D}}_{k}=\frac{\boldsymbol{D}_{k}}{1-\eta^{k+1}}.$ (43)
The derived expression differs from (16) by the shifted exponent. This shift accounts for the fact that the initialization matrix $\boldsymbol{D}_{0}$ should be considered as the zeroth local Hessian estimate $\tilde{\boldsymbol{B}}_{0}$. Lastly, since $0\leq\eta<1$, the denominator in (43) converges to 1 such that
$\hat{\boldsymbol{D}}_{\infty}=\lim_{k\rightarrow\infty}\hat{\boldsymbol{D}}_{k}=\frac{\lim_{k\rightarrow\infty}\boldsymbol{D}_{k}}{\lim_{k\rightarrow\infty}1-\eta^{k+1}}=\boldsymbol{D}_{\infty}.$ (44)
We show empirically in Section 5 that the proposed bias correction improves the Hessian estimate in the early stages and that the limit (44) is also observed in practice.
4 Analytical case study
We first show on an analytical example that the proposed Gauss-Newton approximation yields superlinear convergence in policy optimization. Let us consider the following linear uncertain system
$s^{+}=s+a+w$ (45)
where $w\sim\mathcal{N}(0,\sigma_{w}^{2})$ is a normal distributed uncertainty. The system is controlled with the parameterized policy
$\pi_{\theta}(s)=-\theta^{2}s,$ (46)
where the coefficient is chosen as $\theta^{2}$ instead of $\theta$ to ensure that the approximate Hessian (20) and the Gauss-Newton approximation yield different results due to $\nabla_{\theta}^{2}\pi_{\theta}(s)\neq 0$. The first and second-order policy derivatives are
$\nabla_{\theta}\pi_{\theta}(s)=-2\theta s,\quad\nabla_{\theta}^{2}\pi_{\theta}(s)=-2s.$ (47)
We consider the reward function $r(s,a)=-0.5(s^{2}+a^{2})$. One can show that the state-value function $V^{\pi_{\theta}}(s)$ for the closed-loop system with $0<\gamma<1$ has the following structure
$V^{\pi_{\theta}}(s)=p_{\theta}s^{2}+q_{\theta}$ (48)
where $p_{\theta}$ and $q_{\theta}$ are coefficients that depend on $\theta$. From the definition of state-value function (3) it follows that
$V^{\pi_{\theta}}(s)=$ $-0.5\left(s^{2}+(-\theta^{2}s)^{2}\right)$
$+\gamma\mathbb{E}_{w\sim\mathcal{N}(0,\sigma_{w}^{2})}\left[V^{\pi_{\theta}}(s-\theta^{2}s+w)\right]$
$=$ $\left[-0.5\left(1+\theta^{4}\right)+\gamma\left(1-\theta^{2}\right)^{2}p_{\theta}\right]s^{2}$
$+\gamma\left(q_{\theta}+p_{\theta}\sigma_{w}^{2}\right).$ (49)
From a comparison of coefficients, the following expressions for $p_{\theta}$ and $q_{\theta}$ can be derived
$p_{\theta}=-\frac{0.5\left(1+\theta^{4}\right)}{1-\gamma\left(1-\theta^{2}\right)^{2}},\quad q_{\theta}=\frac{\gamma}{1-\gamma}\sigma_{w}^{2}p_{\theta}.$ (50)
The action-value function $Q^{\pi_{\theta}}(s,a)$ can be written as
$Q^{\pi_{\theta}}(s,a)=$ $-0.5\left(s^{2}+a^{2}\right)$
$+\gamma\mathbb{E}_{w\sim\mathcal{N}(0,\sigma_{w}^{2})}\left[V^{\pi_{\theta}}(s+a+w)\right]$ (51a)
$=$ $-0.5\left(s^{2}+a^{2}\right)+\gamma p_{\theta}(s+a)^{2}$
$+\gamma\left(q_{\theta}+p_{\theta}\sigma_{w}^{2}\right)$ (51b)
with gradient and Hessian with respect to $a$ given as
$\nabla_{a}Q^{\pi_{\theta}}(s,a)=$ $(2\gamma p_{\theta}-1)a+2\gamma p_{\theta}s,$ (52a)
$\nabla_{a}^{2}Q^{\pi_{\theta}}(s,a)=$ $2\gamma p_{\theta}-1$ (52b)
Let the initial state be normally distributed $s_{0}\sim\mathcal{N}(0,\sigma_{0}^{2})$, then the expected cumulative reward $J(\theta)$ can be derived from its definition (5) as
$J(\theta)=\mathbb{E}_{s_{0}\sim\mathcal{N}(0,\sigma_{0}^{2})}\left[V^{\pi_{\theta}}(s_{0})\right]=q_{\theta}+p_{\theta}\sigma_{0}^{2}.$ (53)
The exact gradient and the Hessian with respect to $\theta$ can therefore be derived as
$\nabla_{\theta}J(\theta)=$ $\nabla_{\theta}q_{\theta}+\sigma_{0}^{2}\nabla_{\theta}p_{\theta},$ (54a)
$\nabla_{\theta}^{2}J(\theta)=$ $\nabla^{2}_{\theta}q_{\theta}+\sigma_{0}^{2}\nabla^{2}_{\theta}p_{\theta}.$ (54b)
Setting (54a) to zero and substituting (50) allows to derive the optimal parameter $\theta^{\star}(\gamma)$ as
$\theta^{\star}(\gamma)=\sqrt{\frac{\sqrt{1+4\gamma^{2}}-1}{2\gamma}},$ (55)
which computes to $\theta^{\star}\approx 0.767$ for $\gamma=0.9$.
Next, we will derive the expressions for the Gauss-Newton approximation and the Hessian approximation according to (20). Since the policy gradient $\nabla_{\theta}J(\theta)$ can also be obtained from (12) we have
$\nabla_{\theta}J(\theta)=$ $\mathbb{E}_{s}\left[(-2\theta s)\cdot\left((2\gamma p_{\theta}-1)(-\theta^{2}s)+2\gamma p_{\theta}s\right)\right]$
$=$ $\left(2\theta^{3}(2\gamma p_{\theta}-1)-4\gamma p_{\theta}\theta\right)\mathbb{E}_{s}\left[s^{2}\right],$ (56)
which gives the following expression for $\mathbb{E}_{s}\left[s^{2}\right]$ after equating (54a) and (56)
$\mathbb{E}_{s}\left[s^{2}\right]=\frac{\nabla_{\theta}q_{\theta}+\sigma_{0}^{2}\nabla_{\theta}p_{\theta}}{2\theta^{3}(2\gamma p_{\theta}-1)-4\gamma p_{\theta}\theta}.$ (57)
We derive the matrices $M_{1}(\theta)$ and $M_{2}(\theta)$ from the definitions (21a) and (21b) as
$M_{1}(\theta)=$ $2\theta^{2}(2\gamma p_{\theta}-1)\mathbb{E}_{s}\left[s^{2}\right]-4\gamma p_{\theta}\mathbb{E}_{s}\left[s^{2}\right]$ (58a)
$M_{2}(\theta)=$ $4\theta^{2}(2\gamma p_{\theta}-1)\mathbb{E}_{s}\left[s^{2}\right].$ (58b)
Figure 2 shows the exact Hessian (54b), the approximate Hessian as the sum of (58a) and (58b), and the proposed Gauss-Newton approximation (58b) for a discount factor of $\gamma=0.9$ under variation of $\theta$. All curves intersect at $\theta^{\star}$. Further, the proposed Gauss-Newton approximation remains negative definite over the considered range whereas the more elaborate approximation becomes positive definite far from the optimal parameter.
Figure 2: Hessian and its approximations over $\theta$ for $\gamma=0.9$ and $\sigma_{w}^{2}=\sigma_{0}^{2}=0.1$. The Hessian at $\theta^{\star}$ is highlighted as a star.
Figure 3: Error between the current iterate and the optimal parameter $|\theta_{k}-\theta^{\star}|$ for the second-order approaches and first-order gradient ascent starting at the initial guess $\theta_{0}=0.6$.
Lastly, we investigate the error dynamics $|\theta_{k}-\theta^{\star}|$ between the current iterate $\theta_{k}$ and the optimal parameter $\theta^{\star}$ for the considered second-order approaches and gradient ascent. Figure 3 shows the error dynamics for an initial guess of $\theta_{0}=0.6$. All second-order approaches use a learning rate of $\alpha=1$ whereas the gradient ascent approach uses a learning rate of $\alpha=0.1$. The second-order approaches show a superlinear convergence rate. The first-order gradient ascent algorithm only achieves a linear error reduction.
5 Continuously stirred tank reactor
We demonstrate the improved convergence behavior and data-efficiency of the proposed second-order Gauss-Newton approach in comparison with the state-of-the-art first-order optimizer Adam (Kingma and Ba, 2015), and show that the convergence behavior is similar to more elaborate but more costly approximate Hessian methods like (20) at the example of a continuously stirred tank reactor (CSTR). We also investigate the effect of the introduced momentum for the Hessian and derive tuning rules that we validate with empirical investigations. All code is available online [^1] https://github.com/DeanBrandner/GaussNewtonRLforMPC.
5.1 System representation
The investigated CSTR is taken from Klatt and Engell (1998). We consider the following reaction scheme
$\mathrm{A}\stackrel{{\scriptstyle k_{1}}}{{\longrightarrow}}\mathrm{B}\stackrel{{\scriptstyle k_{2}}}{{\longrightarrow}}\mathrm{C},\quad 2\mathrm{A}\stackrel{{\scriptstyle k_{3}}}{{\longrightarrow}}\mathrm{D}$ (59)
of educt $\mathrm{A}$ to value product $\mathrm{B}$ with the consecutive reaction of $\mathrm{B}$ to side product $\mathrm{C}$ and the parallel reaction of $\mathrm{A}$ to side product $\mathrm{D}$. The reaction constants $k_{i}$ for each reaction are placed above the respective arrows. Figure 4 illustrates the streams entering and leaving the system. A stream with a dilution rate of $F$, a concentration of A $C_{\mathrm{A,in}}$, and a temperature of $T_{\mathrm{in}}$ enters the reactor, and leaves it with the same dilution rate $F$ but concentrations $C_{\mathrm{A}}$, $C_{\mathrm{B}}$, and temperature $T_{\mathrm{R}}$. The CSTR consists of a cooling jacket with an entering heat stream of $\dot{Q}_{\mathrm{in}}$ and a leaving heat stream of $\dot{Q}_{\mathrm{out}}$ with temperature $T_{\mathrm{K}}$. The difference between the entering and leaving heat streams is denoted as $\dot{Q}=\dot{Q}_{\mathrm{in}}-\dot{Q}_{\mathrm{out}}$ and forms together with the dilution rate the manipulated variables $\boldsymbol{a}^{\top}=\boldsymbol{u}^{\top}=(F,\dot{Q})$. The physical state of the system is formed as $\boldsymbol{s}_{\mathrm{phys}}^{\top}=\boldsymbol{x}^{\top}=(C_{\mathrm{A}},C_{\mathrm{B}},T_{\mathrm{R}},T_{\mathrm{K}})$. It is governed by the following ODEs
$\dot{C}_{\mathrm{A}}$ $=F(C_{\mathrm{A,in}}-C_{\mathrm{A}})-k_{1}C_{\mathrm{A}}-k_{3}C_{\mathrm{A}}^{2},$ (60a)
$\dot{C}_{\mathrm{B}}$ $=-FC_{\mathrm{B}}+k_{1}C_{\mathrm{A}}-k_{2}\,C_{\mathrm{B}},$ (60b)
$\dot{T}_{\mathrm{R}}$ $=\frac{k_{1}C_{\mathrm{A}}H_{\mathrm{R,ab}}+k_{2}C_{\mathrm{B}}H_{\mathrm{R,bc}}+k_{3}C_{\mathrm{A}}^{2}H_{\mathrm{R,ad}}}{-\varrho c_{p}}$
$+F(T_{\mathrm{in}}-T_{\mathrm{R}})+\frac{K_{w}\,A_{\mathrm{R}}\,(T_{\mathrm{K}}-T_{\mathrm{R}})}{\varrho\,c_{p}\,V_{\mathrm{R}}},$ (60c)
$\dot{T}_{\mathrm{K}}$ $=\frac{\dot{Q}+K_{w}\,A_{\mathrm{R}}\left(T_{\mathrm{R}}-T_{\mathrm{K}}\right)}{m_{k}c_{p,\mathrm{K}}},$ (60d)
where the reaction rates are defines as
$k_{1}$ $=\mu_{\beta}k_{0,\mathrm{ab}}\exp\left(\frac{-E_{\mathrm{A},\mathrm{ab}}}{T_{\mathrm{R}}+273.15}\right),$ (61a)
$k_{2}$ $=k_{0,\mathrm{bc}}\exp\left(\frac{-E_{\mathrm{A},\mathrm{bc}}}{T_{\mathrm{R}}+273.15}\right),$ (61b)
$k_{3}$ $=k_{0,\mathrm{ad}}\exp\left(\frac{-\mu_{\alpha}E_{\mathrm{A},\mathrm{ad}}}{T_{\mathrm{R}}+273.15}\right).$ (61c)
The parameters are defined as the nominal values in Klatt and Engell (1998). The coefficients $\mu_{\alpha}$ and $\mu_{\beta}$ will be uncertain in the following investigations.
Figure 4: CSTR with entering and leaving streams.
5.2 Agent-environment interaction
The goal is to control the CSTR to a reactor temperature of $T_{\mathrm{R,ref}}=126\,^{\circ}\mathrm{C}$ and a jacket temperature of $T_{\mathrm{K,ref}}=120\,^{\circ}\mathrm{C}$. Both setpoints are obtained by robust steady state optimization. Further, the variations in control input should be kept low to achieve smooth control trajectories. To account for that, the following algebraic equations for the previous control inputs must be introduced
$F_{\mathrm{prev}}=F,\quad\dot{Q}_{\mathrm{prev}}=\dot{Q}.$ (62)
They form the state of previous control inputs as $\boldsymbol{s}_{\mathrm{prev}}^{\top}=(F_{\mathrm{prev}},\dot{Q}_{\mathrm{prev}})$ and complete the definition of the RL state $\boldsymbol{s}^{\top}=(\boldsymbol{s}_{\mathrm{phys}}^{\top},\boldsymbol{s}_{\mathrm{prev}}^{\top})$. The states and actions also need to satisfy constraints  $\boldsymbol{s}_{\mathrm{lb}}\leq\boldsymbol{s}\leq\boldsymbol{s}_{\mathrm{ub}}$ and $\boldsymbol{a}_{\mathrm{lb}}\leq\boldsymbol{a}\leq\boldsymbol{a}_{\mathrm{ub}}$. The upper and lower bounds are listed in Table 1. We define the reward $R(\boldsymbol{s},\boldsymbol{a})$ that encodes the objective above as
$R(\boldsymbol{s},\boldsymbol{a})=$ $R_{\mathrm{ref}}(\boldsymbol{s},\boldsymbol{a})+R_{\Delta\boldsymbol{a}}(\boldsymbol{s},\boldsymbol{a})+R_{\boldsymbol{\sigma}}(\boldsymbol{s},\boldsymbol{a})$ (63a)
$R_{\mathrm{ref}}(\boldsymbol{s},\boldsymbol{a})=$ $-10^{-2}\cdot(T_{\mathrm{R}}-T_{\mathrm{R,ref}})^{2}$
$-10^{-2}\cdot(T_{\mathrm{K}}-T_{\mathrm{K,ref}})^{2}$ (63b)
$R_{\Delta\boldsymbol{a}}(\boldsymbol{s},\boldsymbol{a})=$ $-10^{-2}\cdot\left(\frac{F_{\mathrm{prev}}-F}{F_{\mathrm{max}}-F_{\mathrm{min}}}\right)^{2}$
$-10^{-2}\cdot\left(\frac{\dot{Q}_{\mathrm{prev}}-\dot{Q}}{\dot{Q}_{\mathrm{max}}-\dot{Q}_{\mathrm{min}}}\right)^{2}$ (63c)
$R_{\boldsymbol{\sigma}}(\boldsymbol{s},\boldsymbol{a})=$ $-100\max\left\{0,\boldsymbol{s}_{\mathrm{lb}}-\boldsymbol{s},\boldsymbol{s}-\boldsymbol{s}_{\mathrm{ub}}\right\}$ (63d)
where $R_{\mathrm{ref}}(\boldsymbol{s},\boldsymbol{a})$ encodes the tracking of the references, $R_{\Delta\boldsymbol{a}}(\boldsymbol{s},\boldsymbol{a})$ encodes the low variations in the actions, and $R_{\boldsymbol{\sigma}}(\boldsymbol{s},\boldsymbol{a})$ encodes the constraints.
Table 1: Upper and lower bounds (ub, lb) for states $\boldsymbol{s}$ (left) and actions $\boldsymbol{a}$ (right). Concentrations $C_{i}$ are given in $\mathrm{mol}\,\mathrm{L}^{-1}$, temperatures $T_{i}$ in ${}^{\circ}\mathrm{C}$, dilution rates $F_{i}$ in $\mathrm{h}^{-1}$, and heat flows $\dot{Q}_{i}$ in $\mathrm{kW}$.
$C_{\mathrm{A}}$ $C_{\mathrm{B}}$ $T_{\mathrm{R}}$ $T_{\mathrm{K}}$ $F_{\mathrm{prev}}$ $\dot{Q}_{\mathrm{prev}}$ $F$ $\dot{Q}$
lb 0.10.10.1 0.10.10.1 808080 808080 555 8.5-8.5- 8.5 555 8.5-8.5- 8.5
ub 222 222 140140140 140140140 404040 0 404040 0
The initial states $\boldsymbol{s}_{0}$ of the environment from which all episodes start are sampled uniformly from the feasible state space, so $\boldsymbol{s}_{0}\sim\mathcal{U}(\boldsymbol{s}_{\mathrm{lb}},\boldsymbol{s}_{\mathrm{ub}})$. Further, it is assumed that the coefficients $\mu_{\alpha}$ and $\mu_{\beta}$ are uncertain. Both coefficients are sampled from the uniform distributions $\mu_{\alpha}\sim\mathcal{U}(0.95,1.05)$ and  $\mu_{\beta}\sim\mathcal{U}(0.95,1.05)$ at the beginning of each episode. The sampled realizations of $\mu_{\alpha}$ and $\mu_{\beta}$ are unknown to the agent, which means that the agent needs to learn a control policy that maximizes the cumulative reward but remains robust simultaneously.
The considered agent is a nominal MPC agent that uses the same model as described in (60) and (61), respects the constraints from Table 1, takes the negative reward from (63a) as stage cost, and predicts $N=20$ steps into the future. It is assumed that the parameters are known except for $\mu_{\alpha}$ and $\mu_{\beta}$ as they are uncertain. To improve the MPC policy, we choose to parameterize those coefficients as $\tilde{\mu}_{\alpha}$ and $\tilde{\mu}_{\beta}$ such that
$\tilde{\mu}_{\alpha}=(1+\theta_{\alpha})\cdot 0.95,\quad\tilde{\mu}_{\beta}=(1+\theta_{\beta})\cdot 1.$ (64)
The agent then needs to learn the parameters $\boldsymbol{\theta}^{\top}=(\theta_{\alpha},\theta_{\beta})$ such that the agent performs robust and not too conservative.
We approximate the action-value function using a NN. The following four steps are repeated for each update step in order to ensure a highly accurate Q-function approximation.
  1. 1.
    Sampling of $n_{\mathrm{IC}}$ initial states $\boldsymbol{s}_{0}$ and uncertain coefficient combinations $\mu_{\alpha},\mu_{\beta}$.
  2. 2.
    MPC closed-loop operation with current RL parameters $\boldsymbol{\theta}_{k}$ without exploration for all $n_{\mathrm{IC}}$ initial states, to approximate the state-value function $V_{\boldsymbol{\phi}}(\boldsymbol{s})\approx V^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s})$.
  3. 3.
    Explore for all transitioned states $\boldsymbol{s}$ with exploratory action $\boldsymbol{a}_{\mathrm{exp}}=\pi_{\boldsymbol{\theta}}(\boldsymbol{s})+\mathcal{N}(0,\boldsymbol{\Sigma}_{\mathrm{exp}})$ and based on this approximate $Q_{\boldsymbol{\psi}}(\boldsymbol{s},\boldsymbol{a})\approx Q^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s},\boldsymbol{a})$ locally around the current policy using the state-value function $V_{\boldsymbol{\phi}}(\boldsymbol{s})$.
  4. 4.
    Compute and apply the update according to the desired update rule.
All main training routines are based on the python toolbox do-mpc (Fiedler et al., 2023), which uses a CasADi (Andersson et al., 2019) backend and relies on Ipopt (Wächter and Biegler, 2006).
5.3 Convergence behavior and data-efficiency
We first investigate how our proposed Gauss-Newton approximation (21b) performs compared to the more elaborate Hessian approximation (20) and how it performs compared to the state-of-the-art first-order approach Adam (Kingma and Ba, 2015) with respect to convergence speed and data-efficiency. For all investigated scenarios below, both second-order approaches share the same learning rate of $\alpha=0.1$, gradient momentum of $\beta=0.75$, Hessian momentum of $\eta=0.9$, and initial eigenvalue estimate of $\omega^{-1}=10^{2}$. For the first-order optimizer Adam two different learning rates of $\alpha=10^{-3}$ and $\alpha=10^{-2}$ are investigated. The default Adam hyperparameters of $\beta_{1}=0.9$ and $\beta_{2}=0.999$ are used for the first and second-order moments. Please refer to Kingma and Ba (2015) for more details.
The observed parameter trajectories during training for the initial parameter combination of $\theta_{\alpha,0}=\theta_{\beta,0}=0$ are shown in Figure 5.
Figure 5: Trajectories of $\theta_{\alpha}$ (x-axis) and $\theta_{\beta}$ (y-axis) throughout the policy optimization. All trajectories start at the same initial guess of $\theta_{\alpha,0}=\theta_{\beta,0}=0$. The coloring accounts for $J(\theta_{\alpha},\theta_{\beta})$. The markers denote each considered algorithm. Approx. Newton refers to (20). $\alpha$ denotes the respective learning rates.
Both second-order approaches show stable and fast convergence whereas both Adam approaches struggle. While Adam with a sufficiently small learning rate of $\alpha=10^{-3}$ shows stable but slow training progress, Adam with a larger learning rate of $\alpha=10^{-2}$ appears to have large but unstable improvement throughout the training process. We account the improved convergence behavior of both second-order approaches to the fact that $J(\theta_{\alpha},\theta_{\beta})$ has a significant different curvature in the $\theta_{\alpha}$ and $\theta_{\beta}$-dimensions, which can also be observed from the scales of the $\theta_{\alpha}$ and $\theta_{\beta}$-axis. Contrary to Adam, both second-order approaches can exploit the approximate Hessians to precondition the gradients by their curvature estimate. Adam however must rely on its fixed learning rate, which is limited by the largest element within the gradient. Adam therefore makes rapid progress into the $\theta_{\alpha}$-dimension but struggles along the $\theta_{\beta}$-dimension as the learning rate is limited by $\theta_{\alpha}$. From this, we conclude that second-order approaches are of special interest for MPC as agent for RL since the influence of introduced parameters is difficult to assess beforehand and may vary significantly for the different parameter configurations.
Table 2: Averaged performance metrics of the different training approaches over five initial guesses.
Approach Average time per iteration Number of iterations until convergence Total wall-clock time until convergence Averaged expected cumulative reward
after convergence
$\bar{t}\leavevmode\nobreak\ \mathrm{[min\,iter^{-1}]}$ $N_{\mathrm{conv}}\leavevmode\nobreak\ \mathrm{[iter]}$ $\bar{t}_{\mathrm{conv}}\leavevmode\nobreak\ \mathrm{[min]}$ $J(\theta_{\alpha},\theta_{\beta})$
Adam 1.45 >100 >145 -73.3
Approx. Newton 1.98 65 129 -62.5
Gauss-Newton (prop.) 1.85 65 120 -63.2
We further investigate the robustness of the proposed approaches with respect to the choice of initial guesses. Therefore, we run the training runs with the same specifications as above but for five randomly sampled initial guesses for $\theta_{\alpha,0}\sim\mathcal{U}(-0.05,0.07)$ and $\theta_{\beta,0}\sim\mathcal{U}(-0.75,0.1)$. Figure 6 shows the averaged expected cumulative reward (lines) and its standard deviation (shaded) for the investigated approaches for the five initial guesses. We omit the results for Adam with $\alpha=10^{-2}$ as it shows unsuccessful and unstable learning behavior. All three approaches start with the same averaged initial accumulated reward and robustly improve over the iterations. As observed in Figure 5, the training for Adam with $\alpha=10^{-3}$ is slower than for both second-order approaches. Even after 100 parameter updates, the performance is still not on the same level as both second-order approaches. Both second-order approaches on the other side make rapid improvement in the beginning. The Gauss-Newton approach appears to converge after roughly 65 iterations, whereas the approximate Newton approach still improves marginally.
Figure 6: Average expected cumulative reward (lines) and standard deviation (shaded) for five initial guesses of $\theta_{\alpha}$ and $\theta_{\beta}$ for the investigated policy optimization methods.
Finally we want to point out that not only the number of iterations can be reduced but also the total wall-clock time. During training we measure the computation time per iteration for each training run of the aforementioned agents. Table 2 lists the average wall-clock time per iteration, the number of iterations until convergence, the total wall-clock time until convergence, and the expected cumulative reward after convergence of the different approaches. Although the average computation time per iteration for the proposed Gauss-Newton approach is larger than for Adam, the Gauss-Newton approach only needs 65 iterations to converge to an average expected cumulative reward of $J(\theta_{\alpha},\theta_{\beta})=-63.2$, whereas Adam only achieves a performance of $J(\theta_{\alpha},\theta_{\beta})=-73.3$ after 100 iterations. This results in a total wall-clock time of 120 min for the Gauss-Newton approach and 145 min for Adam. Even in comparison to the approximate Newton approach, the Gauss-Newton approach obtains a 9 min lower total wall-clock time to achieve a similar performance. We expect these improvements to be even more pronounced for more complex systems and MPC policies with higher nonlinearity, which will be the subject of future research.
5.4 Influence of momentum $\beta$ and $\eta$
After we saw that the convergence behavior of the training process can be improved by the proposed Gauss-Newton approach to a similar performance as the considered approximate Newton approach, we also investigate the effect of the momentum on the training stability and training speed for different combinations of $\beta$ and $\eta$.
Figure 7: Expected cumulative reward for the Gauss-Newton approach for $\alpha=0.1$, $\omega^{-1}=10^{1}$ and differing $\beta$ and $\eta$
Figure 7 shows the evolution of the expected cumulative reward $J(\theta_{\alpha},\theta_{\beta})$ for the proposed Gauss-Newton approach with different combinations of $\beta$ and $\eta$. The learning rate and the initial eigenvalue estimate are kept at $\alpha=0.1$ and $\omega^{-1}=10^{1}$. We observe slow but stable convergence if $\eta$ is chosen large such as $\eta=0.99$ regardless of $\beta$ since the exponential moving average is dominated by old values and therefore also by its initialization matrix $\boldsymbol{D}_{0}$. The influence becomes less for smaller momentum such as $\eta=0.9$. Here, the updates are non-conservative from the very beginning and also remain stable. Even smaller or no momentum such as $\eta=0.5$ or $\eta=0$ can cause instability as it can be seen for $\beta=0.75$ and $\eta=0.5$, which diverges after 20 iterations, or $\beta=0.99$ and $\eta=0$, which caused an update that rendered the MPC’s NLP unsolvable. Therefore, the training was aborted. We conclude that stable training results are obtainable for various combinations of $\beta$ and $\eta$. Overall, the robustness of the training algorithm is insensitive with respect to the exact choice of $\eta$ as long as it is chosen sufficiently large. However, in order to also obtain fast convergence, $\eta$ should be chosen not too large.
Finally, we also study the stabilizing effect on the training of the momentum applied to the Hessian estimates at the example of $\eta=0.9$. Figure 8 shows the evolution of the estimated eigenvalues of the Hessian along the training process for the local estimates $\tilde{\boldsymbol{B}}_{k}$, the biased estimate $\boldsymbol{D}_{k}$, and the bias-corrected estimate $\hat{\boldsymbol{D}}_{k}$. It can be seen that the point estimates $\tilde{\boldsymbol{B}}_{k}$ are noisy and that the exponential moving average smoothens out this noise. This is highly important since especially the second eigenvalue varies over more than two orders of magnitude. Further, the bias correction accounts in both cases for the initialization bias enabling conservative regularized Hessian estimates (large eigenvalues) in the beginning. Along the training progress, $\boldsymbol{D}_{k}$ and $\hat{\boldsymbol{D}}_{k}$ converge to the same trajectory as derived in (44).
Figure 8: Eigenvalues of the local Hessian estimate $\tilde{\boldsymbol{B}}_{k}$, the biased averaged Hessian estimate $\boldsymbol{D}_{k}$, and bias-corrected averaged Hessian estimate $\hat{\boldsymbol{D}}_{k}$ for the Gauss-Newton approach over the training steps for $\eta=0.9$ and $\omega^{-1}=10$.
6 Conclusion and future work
We presented a computationally efficient RL framework for MPC policies that enables computationally low-cost second-order policy optimization. The key contribution is a Gauss-Newton approximation of the deterministic policy Hessian that avoids the computation of second-order policy derivatives while still enabling superlinear convergence. The resulting approximation is well-suited for MPC policies as it reduces the amount of expensive agent-environment interactions drastically due to its improved convergence behavior, while it simultaneously avoids to compute the computationally-expensive second-order NLP sensitivities. Furthermore, we proposed a momentum-based exponentially moving averaging scheme for Hessian estimation that improves robustness against noise in local curvature samples. The proposed method improves stability during training and enables faster updates, which is beneficial for sample efficiency.
We rigorously proved the superlinear convergence behavior of the proposed Gauss-Newton approximation. This was confirmed empirically on an analytical example and a nonlinear CSTR benchmark. Results show that the Gauss-Newton approximation in combination with momentum-based averaging significantly reduces the number of parameter updates compared to state-of-the-art first-order methods. In addition, the proposed approach also reduces the total wall-clock time for training compared to first-order and approximate policy Hessian methods while still maintaining robust learning behavior.
Future work will investigate scalability to higher-dimensional MPC policies, integration with robust MPC, and extensions to constrained Markov decision processes.
[…]